# Universal bounds on optimization of free energy harvesting

### 20 Dec 2023 v3

@@ 2.3.2   Constrained kinetics
$\Lambda_{\kappa}^{NO}:=\Lambda^{NO}\cap\{(\bm{p},{\boldsymbol{J}}):J_{NO}\leq\kappa p_{O},J_{ON}\leq\kappa p_{N}\}.$ (S56)
@@
2.4   Proton-pumping as control
@@
$\Lambda^{M_{1}M_{2}}:=\{(\bm{p},{\boldsymbol{J}}):\bm{p}\geq 0,\bm{1}^{T}\bm{p}=1,{\boldsymbol{J}}\geq 0,J_{ij}=0\text{ if }(i,j)\not\in\{{(M_{1},M_{2})},{(M_{2},M_{1})}\}\}\,.$ (S57)
@@
$R^{c}_{M_{2}M_{1}}=\frac{\kappa_{M_{2}M_{1}}}{1+e^{-\Delta s^{\mathrm{tot}}_{M_{2}M_{1}}}}\qquad\qquad R^{c}_{M_{1}M_{2}}=\frac{\kappa_{M_{1}M_{2}}}{1+e^{-\Delta s^{\mathrm{tot}}_{M_{2}M_{1}}}}$ (S58)
with all other $R^{c}_{ij}=0$. The baseline-and-control dynamics correspond to the actual bacteriorhodopsin system described in Sec. 2.1. See also Fig. S4 (right).
Recall that the proton-pumping transition is the only one that involves the membrane potential $\Delta\psi$ (see (S33)), and that it is no longer part of the baseline. For this reason, the optimization problem (5), which only involves baseline parameters, no longer depends on the choice of $\Delta\psi$. We now proceed to study this problem for different choices of additional constraints on $\left(\bm{p},{\boldsymbol{J}}\right)$. For this reason, in Fig. 1, the value for $\mathscr{G}$ when $M_{1}\leftrightarrow M_{2}$ acts as control is a constant. In this case, it is interesting to compare the optimum harvesting rates with the values of $\dot{\mathcal{G}}^{\mathrm{tot}}$ at in vivo membrane potential versus $\max_{\Delta\psi}\dot{\mathcal{G}}^{\mathrm{tot}}$, the maximum value attained by $\dot{\mathcal{G}}^{\mathrm{tot}}$ across all membrane potentials in Fig. 1 (right). Results are shown in Fig. S8.
@@
Activity constraint: analogously to Sec. 2.3, we now define:
$\Lambda_{a}^{M_{1}M_{2}}:=\Lambda^{M_{1}M_{2}}\cap\{(\bm{p}:{\boldsymbol{J}}):J_{M_{1}M_{2}}+J_{M_{2}M_{1}}\leq a,\}$ (S59)
@@
Thermodynamic affinity constraint: owing to the same reasoning as in Sec. 2.3, this constraint proves physically meaningful if combined with another constraint on the fluxes, such as the activity constraint (S59). In this case, we analogously define the feasibility set:
$\Lambda_{\mathcal{A},a}^{M_{1}M_{2}}:=\Lambda_{a}^{M_{1}M_{2}}\cap\left\{(\bm{p},{\boldsymbol{J}}):J_{M_{1}M_{2}}e^{-\mathcal{A}}\leq J_{M_{2}M_{1}}\leq J_{M_{1}M_{2}}e^{\mathcal{A}}\right\},$ (S60)
@@
Kinetic constraint: analogously to Sec. 2.3, we define:
$\Lambda_{\kappa}^{M_{1}M_{2}}:=\Lambda^{M_{1}M_{2}}\cap\{(\bm{p},{\boldsymbol{J}}):J_{M_{1}M_{2}}\leq\kappa p_{M_{2}},J_{M_{2}M_{1}}\leq\kappa p_{M_{1}}\},$ (S61)
@@
2.5   Entire bacteriorhodopsin system treated as baseline
As a final analysis, instead of looking at the optimizability of existing transitions in the bacteriorhodopsin system, we ask to what extent the harvesting rate can be improved by introducing additional control processes. As an example, this may include an additional enzyme that shifts the cycle’s steady-state by permitting a new transition between distant states (e.g. $L\leftrightarrow N$), and in this way increases the rate of proton pumping.
Figure S9: Depiction of the bacteriorhodopsin cycle as the baseline. Here, allowed control processes correspond to additional transitions that may added to the actual dynamics.
In this case, we treat the entire bacteriorhodopsin system as the baseline (see Fig. S9). For the analysis here, we do not introduce additional constraints on the control distributions or fluxes. Then, solving for (6) is equivalent to asking how much additional power can be gained by the addition of any control processes to the actual bacteriorhodopsin system. As shown in Sec. 1.4, in this case, the objective in (5) is achieved in the limit of fast control (${\boldsymbol{J}}\to\infty$ and $\dot{\Sigma}({\boldsymbol{J}})\to 0$). In this case, solving for the maximum harvesting rate corresponds here to solving the simplified optimization problem (6).
Figure S10: Right: the maximum vs. actual harvesting rate as a function of the membrane potential ($\Delta\psi$). Bottom left: actual steady state $\bm{\pi}$ and optimal distribution $\bm{p}$ at in vivo membrane potential value ($\Delta\psi=120\ \text{mV}$) in log-scale. Top left: optimizability (S62) becomes small near the in vivo membrane potential.
Fig. S9 shows the baseline (actual) harvesting rate $\dot{\mathcal{G}}^{\mathrm{tot}}$ and the maximum harvesting rate $\mathscr{G}$ at varying $\Delta\psi$. Interestingly, both peak at around the in vivo values of potential. We can define optimizability as:
$\gamma:=1-\dot{\mathcal{G}}^{\mathrm{tot}}/\mathscr{G},$ (S62)
which quantifies the relative increase of harvesting rate that can be achieved under any possible control. Note that $\gamma$ reaches value 0 when the baseline is already optimal. In Fig. S9 (top left), we see that optimizability reaches a minimum of about $\gamma\approx 0.5$ at around in vivo values. This suggests that, under realistic conditions, bacteriorhodopsin harvests approximately 50% of the fundamental bound given by $\mathscr{G}$. This is remarkably close to optimality, given that $\mathscr{G}$ is derived using very minimal physical assumptions.
Closed-form solutions of Eq. (6) in different regimes
@@ 2.5   Entire bacteriorhodopsin system treated as baseline
3.1   Linear Response (LR) regime
@@
$\sum_{i,j}R^{b}_{ij}p_{j}\ln p_{i}$ $=\sum_{i,j}R^{b}_{ij}p_{j}\left(\ln p_{i}-\ln\pi^{b}_{i}\right)+\sum_{i,j}R^{b}_{ij}p_{j}\ln\pi^{b}_{i}$ (S63)
$=\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\left(\ln p_{i}-\ln\pi^{b}_{i}\right)+\sum_{i,j}R^{b}_{ij}p_{j}\ln\pi^{b}_{i},$ (S64)
where in the last step we used that $-\sum_{i,j}R^{b}_{ij}\pi^{b}_{j}\ln\pi^{b}_{i}=0$ when ${\bm{\pi}^{b}}$ is the steady-state distribution of the baseline. We focus on the first term in (S64), and use the expansion $\ln x\simeq x-1$ for $x\approx 1$, to derive
$\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\ln\left(\frac{p_{i}}{\pi^{b}_{i}}\right)\approx\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\frac{p_{i}-\pi^{b}_{i}}{\pi^{b}_{i}}=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}R^{b}_{ij}z_{j}z_{i},$ (S65)
@@
$\boldsymbol{z}:=D^{-1}(\bm{p}-{\bm{\pi}^{b}})\text{ , with }\ D_{ij}:=\delta_{ij}\sqrt{\pi^{b}_{i}}.$ (S66)
@@
$\boldsymbol{z}^{T}D^{-1}{\bm{\pi}^{b}}=0\rightarrow\sum_{i}\left(p_{i}-\pi^{b}_{i}\right)=0\iff\sum_{i}p_{i}=\sum_{i}\pi^{b}_{i}=1.$ (S67)
@@
$A_{ij}:=\frac{1}{2}\left(R^{b}_{ij}+R^{b}_{ji}\frac{\pi^{b}_{i}}{\pi^{b}_{j}}\right).$ (S68)
@@
$M:=D^{-1}AD\ \rightarrow\ M_{ij}=(D^{-1}AD)_{ij}=\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}A_{ij}.$ (S69)
@@
Going back to (S65), we note that
@@
$=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}\left(R^{b}_{ij}+\frac{\pi^{b}_{i}}{\pi^{b}_{j}}R^{b}_{ji}\right)z_{j}z_{i}=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}z_{j}z_{i}=\sum_{i,j}M_{i,j}z_{j}z_{i}.$ (S70)
In order to keep track of the second term in (S64), let us first recall the free-energy term in (8) and the pay-off vector $\boldsymbol{\phi}$ which, in terms of the baseline parameters, reads as
@@
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}+{\beta}^{-1}{R^{b}}^{T}\ln{\bm{\pi}^{b}}.$ (S71)
For reasons that will become obvious below, we re-scale (S71) by setting
$\boldsymbol{v}:=\frac{\beta}{2}D\boldsymbol{\phi}^{\text{LR}}.$ (S72)
We combine the above with the definition (S66) and the nonlinear term (S70). We then recast the optimization problem in terms of the variable $\boldsymbol{z}$, which is constrained by (S67). Finally, under the LR regime, (6) is approximated by
$\mathscr{G}\simeq\mathscr{G}_{\text{LR}}:={\bm{\pi}^{b}}^{T}\boldsymbol{\phi}+{\beta}^{-1}\max_{\boldsymbol{z}\in\mathbb{R}^{n}:\boldsymbol{z}^{T}D^{-1}{\bm{\pi}^{b}}=0}\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v}.$ (S73)
Eq. (S73) is a quadratic optimization problem, which can be solved using standard techniques from linear algebra. For convenience we summarize these techniques in Lemma 1 in Sec. 5 of this SM. That theorem implies that
$\mathscr{G}_{\text{LR}}={\bm{\pi}^{b}}^{T}\boldsymbol{\phi}-{\beta}^{-1}\boldsymbol{v}^{T}M^{+}\boldsymbol{v}\qquad\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}-DM^{+}\boldsymbol{v}\,$ (S74)
where $M^{+}=\sum_{\alpha>1}\lambda_{\alpha}^{-1}\boldsymbol{m}^{\alpha}{\boldsymbol{m}^{\alpha}}^{T}$ is the Moore-Penrose pseudo-inverse of $M$. In applying Theorem 1, we used result (S129) and the relation $\bm{p}^{*}={\bm{\pi}^{b}}+D\boldsymbol{z}^{*}$, where $\boldsymbol{z}^{*}$ solves (S73).
We can rewrite (S74) using the eigensystem of $M$,
$\mathscr{G}_{\text{LR}}={\bm{\pi}^{b}}^{T}\boldsymbol{\phi}+{\beta}^{-1}\sum_{\alpha>1}\frac{(\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha})^{2}}{-\lambda_{\alpha}}\qquad\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}+\sum_{\alpha>1}\frac{\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}}{-\lambda_{\alpha}}D\boldsymbol{m}^{\alpha}\,$ (S75)
@@
$\Omega_{\alpha}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}=\frac{1}{2}\left(D\boldsymbol{\phi}^{\text{LR}}\right)^{T}\left(D^{-1}\boldsymbol{u}^{\alpha}\right)=\frac{1}{2}\left(\boldsymbol{\phi}^{\text{LR}}\right)^{T}\boldsymbol{u}^{\alpha},$ (S76)
which can be interpreted as the harvesting amplitude of the eigenmodes of $A$. Plugging into (S74) gives Eq. (9) in the main text,
$\mathscr{G}_{\text{LR}}={\bm{\pi}^{b}}^{T}\boldsymbol{\phi}+{\beta}\sum_{\alpha>1}\frac{|\Omega_{\alpha}|^{2}}{-\lambda_{\alpha}}\qquad\text{and}\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}+\sum_{\alpha>1}\frac{\beta\Omega_{\alpha}}{-\lambda_{\alpha}}\boldsymbol{u}^{\alpha}\,$ (S77)
@@
$\max_{\alpha>1}|\Omega_{\alpha}/\lambda_{\alpha}|\ll\frac{1}{\beta(n-1)\max_{i}\sqrt{\pi^{b}_{i}}}\,.$ (S79)
@@ 3.2   Deterministic (D) regime
$\mathscr{G}$ $=\max_{\bm{p}}\,{\beta}^{-1}\left(R^{b}\bm{p}\right)^{T}\ln\bm{p}+\bm{p}^{T}\boldsymbol{\phi}$ (S80)
$\approx\max_{\bm{p}}\bm{p}^{T}\boldsymbol{\phi}=:\mathscr{G}_{\text{D}},$ (S81)
where $(\ln\bm{p})^{T}:=(\ln p_{1},\ldots,\ln p_{n})$. The maximum in (S81) is achieved by $\left(\bm{p}^{*}_{\mathrm{D}}\right)_{i}=\delta_{ii^{*}}$ with $i^{*}=\arg\max_{i}\phi_{i}$, so
$\mathscr{G}_{\text{D}}=\phi_{i^{*}}=\max_{i}\big{[}\dot{g}^{b}_{i}+\sum_{j}R^{b}_{ji}(f_{j}-f_{i}+{g}^{b}_{ji})\big{]}\,,$ (S82)
@@
$\alpha$ $:=\frac{K}{\beta\mathscr{G}_{\text{D}}}=\frac{\max_{i}(-R^{b}_{ii})}{\beta\max_{i}\big{[}\dot{g}^{b}_{i}+\sum_{j}R^{b}_{ji}(f_{j}-f_{i}+{g}^{b}_{ji})\big{]}}\geq 0$ (S83)
$\gamma$ $:=-\ln\Big{(}\min_{i}\pi^{b}_{i}\Big{)}\geq 0\,.$ (S84)
@@
$\left|\frac{\mathscr{G}}{\mathscr{G}_{\text{D}}}-1\right|\leq\alpha\left(-\ln\alpha+\gamma+1\right)\,$ (S85)
The RHS of (S85) vanishes for $\alpha\to 0$, so the LHS tightens as $\mathscr{G}/\mathscr{G}_{\text{D}}\to 1$. We emphasize that the D approximation only becomes accurate in relative, not additive, terms (i.e., it is not necessarily true that $\mathscr{G}-\mathscr{G}_{\text{D}}\to 0$).
To derive (S85), we first consider an upper bound on $\mathscr{G}$. Observe that for any $\bm{p}$,
$-\dot{S}^{b}(\bm{p})=\sum_{i,j}R^{b}_{ij}p_{j}\ln p_{i}\leq\sum_{i}R^{b}_{ii}p_{i}\ln p_{i}\leq\sum_{i}|R^{b}_{ii}|(-p_{i}\ln p_{i})\leq KS(\bm{p})\leq K\ln n$ (S86)
where we used $R^{b}_{ij}p_{j}\ln p_{i}\leq 0$ for $i\neq j$ and $S(\bm{p})\leq\ln n$. Plugging into (S80) and using that $\mathscr{G}_{\text{D}}\geq\bm{p}^{T}\boldsymbol{\phi}$ for all $\bm{p}$ and $\gamma\geq\ln n$, we then have
$\mathscr{G}\leq K(\ln n)/\beta+\mathscr{G}_{\text{D}}=\mathscr{G}_{\text{D}}\alpha\ln n+\mathscr{G}_{\text{D}}\leq\mathscr{G}_{\text{D}}\alpha\gamma+\mathscr{G}_{\text{D}}\,.$ (S87)
To derive a lower bound on $\mathscr{G}$, define the distribution $p_{i}:=\left(1-\alpha\right)\delta_{ii^{*}}+\alpha\pi^{b}_{i}$, with $\alpha$ as in (S83). Plugging this distribution into the objective in (S80) yields
$\mathscr{G}$ $\geq{\beta}^{-1}\left(R^{b}\bm{p}\right)^{T}\ln\bm{p}+\bm{p}^{T}\boldsymbol{\phi}={\beta}^{-1}(1-\alpha)\sum_{i}(R^{b}_{ii^{*}}\ln p_{i})+\bm{p}^{T}\boldsymbol{\phi},$ (S88)
where we used that $\sum_{j}R^{b}_{ij}\left(\left(1-\alpha\right)\delta_{ji^{*}}+\alpha\pi^{b}_{j}\right)=R^{b}_{ii^{*}}(1-\alpha)$ since $R^{b}{\bm{\pi}^{b}}=0$. Using $R^{b}_{i^{*}i^{*}}\ln p_{i^{*}}\geq 0$, we bound the first term on the right side of (S88) as

### 12 Mar 2024 v4

@@ II.3.2 Constrained kinetics
$\Lambda_{\kappa}^{NO}:=\Lambda^{NO}\cap\{(\bm{p},{\boldsymbol{J}}):J_{NO}\leq\kappa p_{O},J_{ON}\leq\kappa p_{N}\}.$ (S2.56)
@@
II.4 Proton-pumping as control
@@
$\Lambda^{M_{1}M_{2}}:=\{(\bm{p},{\boldsymbol{J}}):\bm{p}\geq 0,\bm{1}^{T}\bm{p}=1,{\boldsymbol{J}}\geq 0,J_{ij}=0\text{ if }(i,j)\not\in\{{(M_{1},M_{2})},{(M_{2},M_{1})}\}\}\,.$ (S2.57)
@@
$R^{c}_{M_{2}M_{1}}=\frac{\kappa_{M_{2}M_{1}}}{1+e^{-\Delta s^{\mathrm{tot}}_{M_{2}M_{1}}}}\qquad\qquad R^{c}_{M_{1}M_{2}}=\frac{\kappa_{M_{1}M_{2}}}{1+e^{-\Delta s^{\mathrm{tot}}_{M_{2}M_{1}}}}$ (S2.58)
with all other $R^{c}_{ij}=0$. The baseline-and-control dynamics correspond to the actual bacteriorhodopsin system described in Sec. II.1. See also Fig. S8 (right).
Recall that the proton-pumping transition is the only one that involves the membrane potential $\Delta\psi$ (see (S2.33)), and that it is no longer part of the baseline. For this reason, the optimization problem (10), which only involves baseline parameters, no longer depends on the choice of $\Delta\psi$. We now proceed to study this problem for different choices of additional constraints on $\left(\bm{p},{\boldsymbol{J}}\right)$. For this reason, in Fig. 2, the value for $\mathscr{G}$ when $M_{1}\leftrightarrow M_{2}$ acts as control is a constant. In this case, it is interesting to compare the optimum harvesting rates with the values of $\dot{\mathcal{G}}^{\mathrm{tot}}$ at in vivo membrane potential versus $\max_{\Delta\psi}\dot{\mathcal{G}}^{\mathrm{tot}}$, the maximum value attained by $\dot{\mathcal{G}}^{\mathrm{tot}}$ across all membrane potentials in Fig. 2 (right). Results are shown in Fig. S12.
@@
Activity constraint: analogously to Sec. II.3, we now define:
$\Lambda_{a}^{M_{1}M_{2}}:=\Lambda^{M_{1}M_{2}}\cap\{(\bm{p}:{\boldsymbol{J}}):J_{M_{1}M_{2}}+J_{M_{2}M_{1}}\leq a,\}$ (S2.59)
@@
Thermodynamic affinity constraint: owing to the same reasoning as in Sec. II.3, this constraint proves physically meaningful if combined with another constraint on the fluxes, such as the activity constraint (S2.59). In this case, we analogously define the feasibility set:
$\Lambda_{\mathcal{A},a}^{M_{1}M_{2}}:=\Lambda_{a}^{M_{1}M_{2}}\cap\left\{(\bm{p},{\boldsymbol{J}}):J_{M_{1}M_{2}}e^{-\mathcal{A}}\leq J_{M_{2}M_{1}}\leq J_{M_{1}M_{2}}e^{\mathcal{A}}\right\},$ (S2.60)
@@
Kinetic constraint: analogously to Sec. II.3, we define:
$\Lambda_{\kappa}^{M_{1}M_{2}}:=\Lambda^{M_{1}M_{2}}\cap\{(\bm{p},{\boldsymbol{J}}):J_{M_{1}M_{2}}\leq\kappa p_{M_{2}},J_{M_{2}M_{1}}\leq\kappa p_{M_{1}}\},$ (S2.61)
@@
III Closed-form solutions of Eq. (11) in different regimes
@@
III.1 Linear Response (LR) regime, Eq. (15)
@@
$\sum_{i,j}R^{b}_{ij}p_{j}\ln p_{i}$ $=\sum_{i,j}R^{b}_{ij}p_{j}\left(\ln p_{i}-\ln\pi^{b}_{i}\right)+\sum_{i,j}R^{b}_{ij}p_{j}\ln\pi^{b}_{i}$ (S3.62)
$=\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\left(\ln p_{i}-\ln\pi^{b}_{i}\right)+\sum_{i,j}R^{b}_{ij}p_{j}\ln\pi^{b}_{i},$ (S3.63)
where in the last step we used that $-\sum_{i,j}R^{b}_{ij}\pi^{b}_{j}\ln\pi^{b}_{i}=0$ when ${\bm{\pi}^{b}}$ is the steady-state distribution of the baseline. We focus on the first term in (S3.63), and use the expansion $\ln x\simeq x-1$ for $x\approx 1$, to derive
$\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\ln\left(\frac{p_{i}}{\pi^{b}_{i}}\right)\approx\sum_{i,j}R^{b}_{ij}(p_{j}-\pi^{b}_{j})\frac{p_{i}-\pi^{b}_{i}}{\pi^{b}_{i}}=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}R^{b}_{ij}z_{j}z_{i},$ (S3.64)
@@
$\boldsymbol{z}:=D^{-1}(\bm{p}-{\bm{\pi}^{b}})\text{ , with }\ D_{ij}:=\delta_{ij}\sqrt{\pi^{b}_{i}}.$ (S3.65)
@@
$\boldsymbol{z}^{T}D^{-1}{\bm{\pi}^{b}}=0\rightarrow\sum_{i}\left(p_{i}-\pi^{b}_{i}\right)=0\iff\sum_{i}p_{i}=\sum_{i}\pi^{b}_{i}=1.$ (S3.66)
@@
$A_{ij}:=\frac{1}{2}\left(R^{b}_{ij}+R^{b}_{ji}\frac{\pi^{b}_{i}}{\pi^{b}_{j}}\right).$ (S3.67)
@@
$M:=D^{-1}AD\ \rightarrow\ M_{ij}=(D^{-1}AD)_{ij}=\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}A_{ij}.$ (S3.68)
@@
Going back to (S3.64), we note that
@@
$=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}\left(R^{b}_{ij}+\frac{\pi^{b}_{i}}{\pi^{b}_{j}}R^{b}_{ji}\right)z_{j}z_{i}=\sum_{i,j}\sqrt{\frac{\pi^{b}_{j}}{\pi^{b}_{i}}}z_{j}z_{i}=\sum_{i,j}M_{i,j}z_{j}z_{i}.$ (S3.69)
In order to keep track of the second term in (S3.63), let us first recall the free-energy term in (13) and the pay-off vector $\boldsymbol{\phi}$ which, in terms of the baseline parameters, reads as
@@
$\boldsymbol{\phi}^{\text{LR}}=\boldsymbol{\phi}+{\beta}^{-1}{R^{b}}^{T}\ln{\bm{\pi}^{b}}.$ (S3.70)
For reasons that will become obvious below, we re-scale (S3.70) by setting
$\boldsymbol{v}:=\frac{\beta}{2}D\boldsymbol{\phi}^{\text{LR}}.$ (S3.71)
We combine the above with the definition (S3.65) and the nonlinear term (S3.69). We then recast the optimization problem in terms of the variable $\boldsymbol{z}$, which is constrained by (S3.66). Finally, under the LR regime, (11) is approximated by
$\mathscr{G}\simeq\mathscr{G}_{\text{LR}}:=\boldsymbol{\phi}^{T}{\bm{\pi}^{b}}+{\beta}^{-1}\max_{\boldsymbol{z}\in\mathbb{R}^{n}:\boldsymbol{z}^{T}D^{-1}{\bm{\pi}^{b}}=0}\boldsymbol{z}^{T}M\boldsymbol{z}+2\boldsymbol{z}^{T}\boldsymbol{v}.$ (S3.72)
Eq. (S3.72) is a quadratic optimization problem, which can be solved using standard techniques from linear algebra. For convenience we summarize these techniques in Lemma 1 in Sec. V of this SM. That theorem implies that
$\mathscr{G}_{\text{LR}}=\boldsymbol{\phi}^{T}{\bm{\pi}^{b}}-{\beta}^{-1}\boldsymbol{v}^{T}M^{+}\boldsymbol{v}\qquad\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}-DM^{+}\boldsymbol{v}\,$ (S3.73)
where $M^{+}=\sum_{\alpha>1}\lambda_{\alpha}^{-1}\boldsymbol{m}^{\alpha}{\boldsymbol{m}^{\alpha}}^{T}$ is the Moore-Penrose pseudo-inverse of $M$. In applying Theorem 1, we used result (S5.128) and the relation $\bm{p}^{*}={\bm{\pi}^{b}}+D\boldsymbol{z}^{*}$, where $\boldsymbol{z}^{*}$ solves (S3.72).
We can rewrite (S3.73) using the eigensystem of $M$,
$\mathscr{G}_{\text{LR}}=\boldsymbol{\phi}^{T}{\bm{\pi}^{b}}+{\beta}^{-1}\sum_{\alpha>1}\frac{(\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha})^{2}}{-\lambda_{\alpha}}\qquad\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}+\sum_{\alpha>1}\frac{\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}}{-\lambda_{\alpha}}D\boldsymbol{m}^{\alpha}\,$ (S3.74)
@@
$\Omega_{\alpha}={\beta}^{-1}\boldsymbol{v}^{T}\boldsymbol{m}^{\alpha}=\frac{1}{2}\left(D\boldsymbol{\phi}^{\text{LR}}\right)^{T}\left(D^{-1}\boldsymbol{u}^{\alpha}\right)=\frac{1}{2}\left(\boldsymbol{\phi}^{\text{LR}}\right)^{T}\boldsymbol{u}^{\alpha},$ (S3.75)
which can be interpreted as the harvesting amplitude of the eigenmodes of $A$. Plugging into (S3.73) gives Eq. (15) in the main text,
$\mathscr{G}_{\text{LR}}=\boldsymbol{\phi}^{T}{\bm{\pi}^{b}}+{\beta}\sum_{\alpha>1}\frac{|\Omega_{\alpha}|^{2}}{-\lambda_{\alpha}}\qquad\text{and}\qquad\bm{p}^{*}_{\mathrm{LR}}={\bm{\pi}^{b}}+\sum_{\alpha>1}\frac{\beta\Omega_{\alpha}}{-\lambda_{\alpha}}\boldsymbol{u}^{\alpha}\,$ (S3.76)
@@
$\max_{\alpha>1}|\Omega_{\alpha}/\lambda_{\alpha}|\ll\frac{1}{\beta(n-1)\max_{i}\sqrt{\pi^{b}_{i}}}\,.$ (S3.78)
@@
$\mathscr{G}$ $=\max_{\bm{p}}\,{\beta}^{-1}\left(R^{b}\bm{p}\right)^{T}\ln\bm{p}+\boldsymbol{\phi}^{T}\bm{p}$ (S3.79)
$\approx\max_{\bm{p}}\boldsymbol{\phi}^{T}\bm{p}=:\mathscr{G}_{\text{D}},$ (S3.80)
where $(\ln\bm{p})^{T}:=(\ln p_{1},\ldots,\ln p_{n})$. The maximum in (S3.80) is achieved by $\left(\bm{p}^{*}_{\mathrm{D}}\right)_{i}=\delta_{ii^{*}}$ with $i^{*}=\arg\max_{i}\phi_{i}$, so
$\mathscr{G}_{\text{D}}=\phi_{i^{*}}=\max_{i}\big{[}\dot{g}^{b}_{i}+\sum_{j}R^{b}_{ji}(f_{j}-f_{i}+{g}^{b}_{ji})\big{]}\,,$ (S3.81)
@@
$\alpha$ $:=\frac{K}{\beta\mathscr{G}_{\text{D}}}=\frac{\max_{i}(-R^{b}_{ii})}{\beta\max_{i}\big{[}\dot{g}^{b}_{i}+\sum_{j}R^{b}_{ji}(f_{j}-f_{i}+{g}^{b}_{ji})\big{]}}\geq 0$ (S3.82)
$\gamma$ $:=-\ln\Big{(}\min_{i}\pi^{b}_{i}\Big{)}\geq 0\,.$ (S3.83)
@@
$\left|\frac{\mathscr{G}}{\mathscr{G}_{\text{D}}}-1\right|\leq\alpha\left(-\ln\alpha+\gamma+1\right)\,$ (S3.84)
The RHS of (S3.84) vanishes for $\alpha\to 0$, so the LHS tightens as $\mathscr{G}/\mathscr{G}_{\text{D}}\to 1$. We emphasize that the D approximation only becomes accurate in relative, not additive, terms (i.e., it is not necessarily true that $\mathscr{G}-\mathscr{G}_{\text{D}}\to 0$).
To derive (S3.84), we first consider an upper bound on $\mathscr{G}$. Observe that for any $\bm{p}$,
$-\dot{S}^{b}(\bm{p})=\sum_{i,j}R^{b}_{ij}p_{j}\ln p_{i}\leq\sum_{i}R^{b}_{ii}p_{i}\ln p_{i}\leq\sum_{i}|R^{b}_{ii}|(-p_{i}\ln p_{i})\leq KS(\bm{p})\leq K\ln n$ (S3.85)
where we used $R^{b}_{ij}p_{j}\ln p_{i}\leq 0$ for $i\neq j$ and $S(\bm{p})\leq\ln n$. Plugging into (S3.79) and using that $\mathscr{G}_{\text{D}}\geq\boldsymbol{\phi}^{T}\bm{p}$ for all $\bm{p}$ and $\gamma\geq\ln n$, we then have
$\mathscr{G}\leq K(\ln n)/\beta+\mathscr{G}_{\text{D}}=\mathscr{G}_{\text{D}}\alpha\ln n+\mathscr{G}_{\text{D}}\leq\mathscr{G}_{\text{D}}\alpha\gamma+\mathscr{G}_{\text{D}}\,.$ (S3.86)
To derive a lower bound on $\mathscr{G}$, define the distribution $p_{i}:=\left(1-\alpha\right)\delta_{ii^{*}}+\alpha\pi^{b}_{i}$, with $\alpha$ as in (S3.82). Plugging this distribution into the objective in (S3.79) yields
$\mathscr{G}$ $\geq{\beta}^{-1}\left(R^{b}\bm{p}\right)^{T}\ln\bm{p}+\boldsymbol{\phi}^{T}\bm{p}={\beta}^{-1}(1-\alpha)\sum_{i}(R^{b}_{ii^{*}}\ln p_{i})+\boldsymbol{\phi}^{T}\bm{p},$ (S3.87)
where we used that $\sum_{j}R^{b}_{ij}\left(\left(1-\alpha\right)\delta_{ji^{*}}+\alpha\pi^{b}_{j}\right)=R^{b}_{ii^{*}}(1-\alpha)$ since $R^{b}{\bm{\pi}^{b}}=0$. Using $R^{b}_{i^{*}i^{*}}\ln p_{i^{*}}\geq 0$, we bound the first term on the right side of (S3.87) as